OpenStudy (anonymous):

Two end points of a line segment are (-12, -2) and (-6, -10). What are the coordinates of the point on the line through which its bisector passes? (-18, -12) (-9, -6) (-7, -8) (-11, -4)

OpenStudy (ghazi):

use mid point formula (x1+x2)/2, (y1+y2)/2

OpenStudy (anonymous):

I've been stuck on this question for the longest time, I really don't know how to do that

OpenStudy (anonymous):

I'm so confused and I'm stuck on these last 5 questions and I don't know what else to do.

OpenStudy (ghazi):

can you do it now?

OpenStudy (anonymous):

@ghazi no.. i have no idea how to get the answer

OpenStudy (ghazi):

(-12-6)/2= -18/2 and (-2-10)/2= -12/2.....now you can simplify it

OpenStudy (anonymous):

-9, -6, thank you